# Zero-shot Benchmarking: A Framework for Flexible and Scalable Automatic Evaluation of Language Models

### v1

@@ Abstract
As language models improve and become capable of performing more complex tasks across modalities, evaluating them automatically becomes increasingly challenging. Developing strong and robust task-specific automatic metrics gets harder, and human-annotated test sets—which are expensive to create—saturate more quickly. A compelling alternative is to design reliable strategies to automate the creation of test data and evaluation, but previous attempts either rely on pre-existing data, or focus solely on individual tasks. We present Zero-shot Benchmarking (ZSB), a framework for creating high-quality benchmarks for any task by leveraging language models for both synthetic test data creation and evaluation. ZSB is simple and flexible: it requires only the creation of a prompt for data generation and one for evaluation; it is scalable to tasks and languages where collecting real-world data is costly or impractical; it is model-agnostic, allowing the creation of increasingly challenging benchmarks as models improve. To assess the effectiveness of our framework, we create benchmarks for five text-only tasks and a multi-modal one: general capabilities in four languages (English, Chinese, French, and Korean), translation, and general vision-language capabilities in English. We then rank a broad range of open and closed systems on our benchmarks. ZSB rankings consistently correlate strongly with human rankings, outperforming widely-adopted standard benchmarks. Through ablations, we find that strong benchmarks can be created with open models, and that judge model size and dataset variety are crucial drivers of performance. We release all our benchmarks, and code to reproduce our experiments and to produce new benchmarks.
@@ 1 Introduction
ZSB represents a shift in how we approach benchmarking, particularly for specialized or emerging tasks. Unlike traditional benchmarks that require extensive human-annotated data, ZSB can be rapidly deployed for applications where collecting real-world data would be too costly or impractical, like evaluating chat capabilities in non-English languages. ZSB is also flexible and modular, allowing users to seamlessly create benchmarks that evaluate several skills at once, like vision-language chat capabilities, which combine textual and visual understanding. Furthermore, contrary to static test sets, the quality of Zero-shot benchmarks evolves alongside improvements in LLM capabilities: as more powerful models emerge, both the data generation and evaluation components can be updated simply by switching to newer models, ensuring benchmarks remain relevant and challenging.
To assess the effectiveness of ZSB, we create benchmarks for five text tasks and a multi-modal one: general capabilities [^1] We define “general capabilities” in Section 3.2. in four languages (English, Chinese, French, and Korean), translation (across eleven language pairs), and vision-language general capabilities in English. We evaluate up to 22 LLMs and VLMs on our benchmarks, yielding rankings that correlate strongly with human rankings, often outperforming widely-adopted standard benchmarks. We also perform a series of ablations to show that good benchmarks can be created with open models, and that judge model size and dataset variety are crucial factors for guaranteeing the best performance. Accompanying this work, we release all our benchmarks, as well as code to reproduce our experiments and to produce new benchmarks. [^2] Datasets and code available on Github.
@@ 2 Zero-shot Benchmarking (ZSB)
Evaluation requires creating a judgment prompt that can contain a reference answer. Contrary to previous works, we opt for a direct assessment (Zheng et al., 2023; Liu et al., 2023; Ye et al., 2023; Kim et al., 2023; Graham et al., 2018; 2015, DA) prompt with a 6-point Likert scale, instead of a pairwise comparison prompt (Zheng et al., 2023; Zhao et al., 2024; Luo et al., 2024, PWC). Though DAs pose the challenge of defining an appropriate scoring scale, they are more versatile than PWC evaluation in that they do not require multiple rounds of comparisons to obtain an Elo ranking (Zheng et al., 2023), or the definition of a baseline answer, as done by Li et al. (2024c). Furthermore, a DA prompt can leverage pre-established scoring systems, such as the MQM framework for translation (Lommel et al., 2014; Burchardt, 2013). We show that both DA and PWC yield system rankings of similar quality (§5). All judgment prompts used are included in Appendix A.3. In this work, we only leverage a single model for evaluation, [^3] We test and compare several different open and closed LLMs as judges (§5). but we note that ZSB can be coupled with human-in-the-loop evaluation, or approaches that attempt to mitigate the biases of a single judge, like querying multiple judges (Verga et al., 2024; Zhao et al., 2024).
@@ 3 Experimental Setup
For each task, we create 500 test instances using Claude-Sonnet-3.5444We choose Claude-Sonnet-3.5 because it is a state-of-the-art LLM, but we try other models (§5). and evaluate up to 22 systems. We choose a wide variety of open and closed systems of different sizes and families; the complete list of evaluated systems per task, as well as their standings in the gold and ZSB rankings, can be found in Appendix B.1. Evaluation is performed with Claude-Sonnet-3.5 using a 6-point Likert scale in the prompt. Systems are ranked according to their average score across all instances.
@@ 3.2 Tasks
The term “general capabilities” is often used to refer to the real-world utility of language models in addressing queries that involve core knowledge, instruction-following, and conversational capabilities (Zheng et al., 2023). However, traditional test sets often rely on multiple-choice assessments, which have inherent limitations, such as lacking diversity and complexity. The Chatbot Arena (Zheng et al., 2023) offers an alternative approach by crowdsourcing open-ended queries and pairwise response evaluations. Its main advantage is that it is closer to a deployment of an LLM: humans will create diverse and complex queries as they would in a real-world setting. Additionally, by relying on human evaluation, its rankings will reflect human preferences more than multiple-choice accuracy. With over 1 million queries and evaluations, it has become the de facto standard for general capabilities evaluation. With ZSB, we want to mimic this approach, but fully automatically. As such, we generate open-ended queries across English, Chinese, French, and Korean and evaluate up to 22 LLMs, comparing our resulting rankings with the Arena’s.
@@ 3.3 Baselines
For LLM general capabilities, we consider an additional baseline: M-ArenaHard (Dang et al., 2024), which was obtained by translating ArenaHard (Li et al., 2024c) with Google Translate. ArenaHardAuto contains 500 challenging Chatbot Arena queries in English. We use our judging framework and not the original implementation to isolate the impact of the test data’s provenance on ranking performance. We consider this to be an upper bound on the performance of our benchmarks, as the queries of ArenaHard were taken directly from the source of the gold rankings we use.
4 Experiments
@@ 4.1 LLM General Capabilities
Table 4 shows the PA scores for each language. ZSB outperforms averaging test sets in standard benchmarks—one of the most used method for evaluating general capabilities—across the board. Furthermore, it outperforms Borda count—the strongest baseline—on French and Korean. The same trend is observed when comparing against almost all individual benchmarks, instead of aggregations (see Appendix B.2). Remarkably, ZSB outperforms the upper bound in French at 0.845 PA. These findings outline the potential of ZSB for creating high-quality multilingual benchmarks at a low cost: our English test set cost roughly \$5, while ArenaHardAuto (Li et al., 2024c) cost \$500, and GPQA (Rein et al., 2023)—which got 0.73 PA—cost upward of \$100,000 (Zheng et al., 2023).
@@ 4.2 Translation
Table 2 shows PA scores for MT, aggregated across language pairs. At 0.79 PA, ZSB is only two points below AutoRank and CometKiwi-XXL, the state-of-the-art reference-free metric for MT. ZSB is also close to reference-based xComet and MetricX. This is a remarkable result given the complexity of the task and the quality of the baseline: not only are the metrics we report tailored for MT, but the data underlying their system rankings is the same as the gold rankings. Because of the high cost of hiring professionals to create data and evaluate systems, WMT lacks language coverage and many flavours of translation (e.g., transcreation, which mixes translation with rewriting for cultural adaptation). ZSB could potentially be used for creating benchmarks for any language and task (or mixture thereof) with minimal human effort, which may alleviate this issue.
@@ 4.3 VLM English General Capabilities
Table 4 shows that ZSB outperforms standard benchmarks by a considerable margin—3 PA points—regardless of the aggregation method. This is a strong result, considering that the leaderboard leverages 8 datasets, while our dataset contains only 500 instances. Notably, all but one test set, MMMU, are outperformed by ZSB (see Table 22 in the Appendix). This highlights the flexibility of Zero-shot Benchmarking in seamlessly extending to multimodal settings.
@@ 5 Ablations
The vast majority benchmarks comprised of synthetic data were generated with proprietary models (Fan et al., 2023; Sprague et al., 2023; Wu et al., 2024). Using open models for data generation is an attractive alternative for three reasons: i) data may be cheaper to generate; ii) the resulting data can be free of usage restrictions; and, perhaps most importantly, iii) it is easier to improve benchmarks as ZSB improves in tandem with open model advancements. Thus, we evaluate the performance of open models Llama-3.3-70B-Instruct (Grattafiori et al., 2024) and Qwen2.5-{3, 7, 14, 32, 72}B-Instruct (Yang et al., 2024b) on both the data generation and judgment components of our framework. Table 3 shows that benchmarks created with Llama-3.3-70B-Instruct or Qwen2.5-72B-Instruct can achieve strong correlations with human rankings in English, and, remarkably, outside of English (see Appendix C.2). Nevertheless, choosing the right model is crucial for the quality of the benchmark—for example, Qwen is especially strong in Korean, while Llama lags behind.
@@ A.1 ZSB Meta-prompts for Data Generation
$<$START OF PROMPT$>$
[[[[Your generated prompt here. IMPORTANT: include only the prompt.]]]]
$<$END OF PROMPT$>$
$<$START OF REFERENCE$>$
[[[[A reference answer to your prompt. IMPORTANT: include only the reference answer.]]]]
$<$END OF REFERENCE$>$
@@ 1. Review the following input variables:
$<$START OF SOURCE$>$
[[[[INSERT THE SOURCE TEXT HERE]]]]
$<$END OF SOURCE$>$
$<$START OF REFERENCE TRANSLATION$>$
[[[[INSERT THE REFERENCE TRANSLATION HERE]]]]
$<$END OF REFERENCE TRANSLATION$>$
@@
$<$START OF PROMPT$>$
[[[[Your generated prompt here. This should be a question or instruction that will be paired with the image for the model to analyze.]]]]
$<$END OF PROMPT$>$
$<$START OF REFERENCE$>$
[[[[A reference answer, and only the answer, to the generated prompt.]]]]
$<$END OF REFERENCE$>$
@@ A.2 Meta-prompt Attributes and Distributions
Table 15: Part 11 of all meta-prompt attribute values and their prevalence on all Zero-Shots Benchmarks released. Distributions are the same across languages and language pairs, respectively. The VLM task meta-prompt does not have attributes, only an image.
@@ A.3 ZSB Judgement Prompts
{“feedback”: “<<<detailed explanation of the score based on the criteria>>>”, “result”: “<<<only a number from 1 to 6>>>”}
@@
$<$END OF PROMPT$>$
$<$START OF ANSWER$>$
@@
{“feedback”: “<<<detailed explanation of the score based on the criteria>>>”, “result”: “<<<only a number from 1 to 6>>>”}
$<$START OF SOURCE TEXT$>$
@@
$<$END OF SOURCE TEXT$>$
$<$START OF TRANSLATION$>$
@@
{“feedback”: “”<<<detailed explanation of the score based on the criteria>>>”, “result”: “<<<only a number from 1 to 6>>>”}
@@
<<<START OF PROMPT>>> \${prompt} <<<END OF PROMPT>>>
$<$START OF ANSWER$>$
@@ A.5 ZSB Safety Metadata
{“feedback”: “<<<detailed explanation of the score based on the criteria>>>”, “result”: “<<<only a number from 1 to 6>>>”}
@@
{“feedback”: “<<<detailed explanation of the score based on the criteria>>>”, “result”: “<<<only a number from 1 to 6>>>”}
@@
$<$END OF PROMPT$>$
$<$START OF SAFETY FEEDBACK$>$
@@
$<$END OF PROMPT$>$
$<$START OF SAFETY FEEDBACK$>$
@@
$<$END OF PROMPT$>$
$<$START OF SAFETY FEEDBACK$>$

### v2

@@ Abstract
As language models improve and grow capable of performing more complex tasks across modalities, evaluating them automatically becomes increasingly challenging. Developing strong and robust task-specific automatic metrics gets harder, and human-annotated test sets—which are expensive to create—saturate more quickly. A compelling alternative is to design reliable strategies to automate the creation of test data and evaluation, but previous attempts either rely on pre-existing data, or focus solely on individual tasks. We present Zero-shot Benchmarking (ZSB), a framework for creating high-quality benchmarks for any task by leveraging language models for both synthetic test data creation and evaluation. ZSB is simple and flexible: it requires only the creation of a prompt for data generation and one for evaluation; it is scalable to tasks and languages where collecting real-world data is costly or impractical; it is model-agnostic, allowing the creation of increasingly challenging benchmarks as models improve. To assess the effectiveness of our framework, we create benchmarks for five text-only tasks and a multi-modal one: general capabilities in four languages (English, Chinese, French, and Korean), translation, and general vision-language capabilities in English. We then rank a broad range of open and closed systems on our benchmarks. ZSB rankings consistently correlate strongly with human rankings, outperforming widely-adopted standard benchmarks. Through ablations, we find that strong benchmarks can be created with open models, and that judge model size and dataset variety are crucial drivers of performance. We release all our benchmarks, and code to reproduce our experiments and to produce new benchmarks. [^1] All data, prompts, and code we used and to create and run new benchmarks is on Github.
@@ 1 Introduction
ZSB represents a shift in how we approach benchmarking, particularly for specialized or emerging tasks. Unlike traditional benchmarks that require extensive human-annotated data, ZSB can be rapidly deployed for applications where collecting real-world data would be too costly or impractical, like evaluating chat capabilities in non-English languages. ZSB is also flexible and modular, allowing users to seamlessly create benchmarks that evaluate several skills at once, like vision-language chat capabilities, which combine textual and visual understanding. Furthermore, contrary to static test sets, and because the framework is model-agnostic, the quality of Zero-shot benchmarks evolves alongside improvements in LLM capabilities: as more powerful models emerge, both the data generation and evaluation components can be updated simply by switching to newer models, ensuring benchmarks remain relevant and challenging.
To assess the effectiveness of ZSB, we create benchmarks for five text tasks and a multi-modal one: general capabilities [^2] We define “general capabilities” in Section 3.2. in four languages (English, Chinese, French, and Korean), translation (across eleven language pairs), and vision-language general capabilities in English. We evaluate up to 22 LLMs and VLMs on our benchmarks, yielding rankings that correlate strongly with human rankings, often outperforming widely-adopted standard benchmarks. We also perform a series of ablations to show that good benchmarks can be created with open models, and that judge model size and dataset variety are crucial factors for guaranteeing the best performance. Accompanying this work, we release all our benchmarks, as well as code to reproduce our experiments and to produce new benchmarks. [^3] Datasets and code will be released upon publication.
@@ 2 Zero-shot Benchmarking (ZSB)
Evaluation requires creating a judgment prompt that can contain a reference answer. Contrary to previous works, we opt for a direct assessment (Zheng et al., 2023; Liu et al., 2023; Ye et al., 2023; Kim et al., 2023; Graham et al., 2018; 2015, DA) prompt with a 6-point Likert scale, instead of a pairwise comparison prompt (Zheng et al., 2023; Zhao et al., 2024; Luo et al., 2024, PWC). Though DAs pose the challenge of defining an appropriate scoring scale, they are more versatile than PWC evaluation in that they do not require multiple rounds of comparisons to obtain an Elo ranking (Zheng et al., 2023), or the definition of a baseline answer, as done by Li et al. (2024c). Furthermore, a DA prompt can leverage pre-established scoring systems, such as the MQM framework for translation (Lommel et al., 2014; 2013). We show that both DA and PWC yield system rankings of similar quality (§5). All judgment prompts used are included in Appendix A.3. In this work, we only leverage a single model for evaluation, [^4] We test and compare several different open and closed LLMs as judges (§5). but we note that ZSB can be coupled with human-in-the-loop evaluation, or approaches that attempt to mitigate the biases of a single judge, like querying multiple judges (Verga et al., 2024; Zhao et al., 2024).
@@ 3 Experimental Setup
For each task, we create 500 test instances using Claude-Sonnet-3.5555We choose Claude-Sonnet-3.5 because it is a state-of-the-art LLM, but we try other models (§5). (totaling 8000 instances) and evaluate up to 22 systems. We choose a wide variety of open and closed systems of different sizes and families; the complete list of evaluated systems per task, as well as their standings in the gold and ZSB rankings, can be found in Appendix B.1. Evaluation is performed with Claude-Sonnet-3.5 using a 6-point Likert scale in the prompt. Systems are ranked according to their average score across all instances.
@@ 3.2 Tasks
The term “general capabilities” is often used to refer to the real-world utility of language models in addressing queries that involve core knowledge, instruction-following, and conversational capabilities (Zheng et al., 2023). However, traditional test sets often rely on multiple-choice assessments, which have inherent limitations, such as lacking diversity and complexity. The Chatbot Arena (Zheng et al., 2023) offers an alternative approach by crowdsourcing open-ended queries and pairwise response evaluations. Its main advantage is that it is closer to a deployment of an LLM: humans will create diverse and complex queries as they would in a real-world setting. Additionally, by relying on human evaluation, its rankings will reflect human preferences more than multiple-choice accuracy. With over 1 million queries and evaluations, it has become the de facto standard for general capabilities evaluation. With ZSB, we want to mimic this approach, but fully automatically. As such, we generate open-ended queries across English, Chinese, French, and Korean and evaluate up to 22 LLMs, comparing our resulting rankings with the Arena’s. We include a brief per-category analysis of the ArenaHard data and our final datasets in Appendix D.
@@ 3.3 Baselines
For LLM general capabilities, we consider an additional baseline: M-ArenaHard (Dang et al., 2024), which was obtained by translating ArenaHard (Li et al., 2024c) with Google Translate. ArenaHard contains 500 challenging Chatbot Arena queries in English. We use our judging framework and not the original implementation to isolate the impact of the test data’s provenance on ranking performance. We consider this to be an upper bound on the performance of our benchmarks, as the queries of ArenaHard were taken directly from the source of the gold rankings we use.
4 Main Results
@@ 4.1 LLM General Capabilities
Table 4 shows the PA scores for each language. ZSB outperforms averaging test sets in standard benchmarks—one of the most used method for evaluating general capabilities—across the board. Furthermore, it outperforms Borda count—the strongest baseline—on French and Korean. The same trend is observed when comparing against almost all individual benchmarks, instead of aggregations (see Appendix B.2). Remarkably, ZSB outperforms the upper bound in French at 0.845 PA. These findings outline the potential of ZSB for creating high-quality multilingual benchmarks at a low cost: our English test set cost roughly \$5, while ArenaHard (Li et al., 2024c) cost \$500, and GPQA (Rein et al., 2023)—which got 0.73 PA—cost upward of \$100,000 (Zheng et al., 2023).
@@ 4.2 Translation
Table 2 shows PA scores for MT, aggregated across language pairs (we show by-LP results in Appendix C.5). At 0.79 PA, ZSB is only two points below AutoRank and CometKiwi-XXL, the state-of-the-art reference-free metric for MT. ZSB is also close to reference-based xComet and MetricX. This is a remarkable result given the complexity of the task and the quality of the baseline: not only are the metrics we report tailored for MT, but the data underlying their system rankings is the same as the gold rankings. Because of the high cost of hiring professionals to create data and evaluate systems, WMT lacks language coverage and many flavours of translation (e.g., transcreation, which mixes translation with rewriting for cultural adaptation). ZSB could potentially be used for creating benchmarks for any language and task (or mixture thereof) with minimal human effort, which may alleviate this issue.
@@ 4.3 VLM English General Capabilities
Table 4 shows that ZSB outperforms standard benchmarks by a considerable margin—4.5 PA points—regardless of the aggregation method. This is a strong result, considering that the leaderboard leverages 8 datasets, while our dataset contains only 500 instances. Notably, all but one test set, MMMU, are outperformed by ZSB (see Table 23 in the Appendix). This highlights the flexibility of Zero-shot Benchmarking in seamlessly extending to multimodal settings.
@@ 5 Ablations
The vast majority benchmarks comprised of synthetic data were generated with proprietary models (Fan et al., 2023; Sprague et al., 2023; Wu et al., 2024). Using open models for data generation is an attractive alternative for three reasons: i) data may be cheaper to generate; ii) the resulting data can be free of usage restrictions; and, perhaps most importantly, iii) it is easier to improve benchmarks as ZSB improves in tandem with open model advancements. Thus, we evaluate the performance of open models Llama-3.3-70B-Instruct (Grattafiori et al., 2024) and Qwen2.5-{3, 7, 14, 32, 72}B-Instruct (Yang et al., 2024b) on both the data generation and judgment components of our framework. Table 3 shows that benchmarks created with Llama-3.3-70B-Instruct or Qwen2.5-72B-Instruct can achieve strong correlations with human rankings in English, and, remarkably, outside of English (see Appendix C.2). Importantly, Qwen outperforms averaging standard benchmarks across the board. Choosing the right model is crucial for the quality of the benchmark—for example, Qwen is especially strong in Korean, while Llama lags behind. Finally, to further illustrate that the quality of ZSB can improve alongside that of open models, we present results for general capabilities using Llama 4 Maverick (Llama 4 Team, 2025) as a data generator and judge in Appendix C.4, showing it outperforms Llama 3.3 on 3 out of 4 languages.
@@
Data instances adhere to meta-prompt specifications, and benchmark performance is robust to meta-prompt variations, but reference answer quality is contentious.
Throughout this work, we only used a single meta-prompt for creating general capabilities data, and we mostly performed reference-less evaluation. However, one step to ensure the practical utility and versatility of our framework is to verify its reliability with respect to the reference answer it provides and to the meta-prompt—both in terms of whether the generated instances respect it, and whether useful instances are still generated after modifying the prompt lexically. Thus, in Table 5 we present five statistics for all the languages we consider: (i) the percentage of data instances that abide by the meta-prompt’s attributes; (ii) the percentage of perfect references (scored 1 to 6); (iii) the percentage of score 5 references; (iv) and (v) the PA average and standard deviation of our benchmarks generated with three meta-prompts (the default prompt plus two rewrites). [^9] We prompt Claude-Sonnet-3.5 to obtain all statistics; the scripts and prompts used are in the Github repository linked at the beginning of this work. The statistics show similar patterns across languages. While the generated data is consistent with meta prompt specifications, reference quality could be improved: the vast majority of references are quite good (score 5), but only a small fraction are perfect.
LLM General Capabilities
English Chinese French Korean
Baselines
% consistency w/ meta prompt attributes 95.8 97.4 99.6 99.0
% perfect references 7.4 26.8 5.2 2.0
% score 5 references 89.2 68.6 86.6 88.2
% Avg. PA for 3 meta prompts 0.8470 0.8341 0.8283 0.8419
% Std. dev. of PA for 3 meta prompts 0.0090 0.0025 0.0175 0.0074
Table 5: Data statistics towards ensuring the practical robustness and versatility of our data generation framework: (i) the percentage of data instances that abide by the meta-prompt’s attributes; (ii) the percentage of perfect references (scored 1 to 6); (iii) the percentage of score 5 references; (iv) and (v) the PA average and standard deviation of our benchmarks generated with three meta-prompts (the default prompt plus two rewrites).
@@ A.1 ZSB Meta-prompts for Data Generation
<<START OF PROMPT>>
[[Your generated prompt here. IMPORTANT: include only the prompt.]]
<<END OF PROMPT>>
<<START OF REFERENCE>>
[[A reference answer to your prompt. IMPORTANT: include only the reference answer.]]
<<END OF REFERENCE>>
@@ 1. Review the following input variables:
<<START OF SOURCE>>
[[INSERT THE SOURCE TEXT HERE]]
<<END OF SOURCE>>
<<START OF REFERENCE TRANSLATION>>
[[INSERT THE REFERENCE TRANSLATION HERE]]
<<END OF REFERENCE TRANSLATION>>
@@
<<START OF PROMPT>>
[[Your generated prompt here. This should be a question or instruction that will be paired with the image for the model to analyze.]]
<<END OF PROMPT>>
<<START OF REFERENCE>>
[[A reference answer, and only the answer, to the generated prompt.]]
<<END OF REFERENCE>>
@@ A.3 ZSB Judgement Prompts
{“feedback”: “<<detailed explanation of the score based on the criteria>>”, “result”: “<<only a number from 1 to 6>>”}
@@
<<END OF PROMPT>>
<<START OF ANSWER>>
@@
{“feedback”: “<<detailed explanation of the score based on the criteria>>”, “result”: “<<only a number from 1 to 6>>”}
<<START OF SOURCE TEXT>>
@@
<<END OF SOURCE TEXT>>
<<START OF TRANSLATION>>
@@
{“feedback”: “”<<detailed explanation of the score based on the criteria>>”, “result”: “<<only a number from 1 to 6>>”}
@@
<<START OF PROMPT>> \${prompt} <<END OF PROMPT>>
<<START OF ANSWER>>
@@ A.5 ZSB Safety Metadata
{“feedback”: “<<detailed explanation of the score based on the criteria>>”, “result”: “<<only a number from 1 to 6>>”}
@@
{“feedback”: “<<detailed explanation of the score based on the criteria>>”, “result”: “<<only a number from 1 to 6>>”}
@@
<<END OF PROMPT>>
<<START OF SAFETY FEEDBACK>>
@@
<<END OF PROMPT>>
<<START OF SAFETY FEEDBACK>>
@@
<<END OF PROMPT>>
<<START OF SAFETY FEEDBACK>>